PS-IS-119/RC

 

MINUTES OF THE PS-ISOLDE COORDINATION MEETING

held on 27 August 2002

 

 

Present:

S. Burger, R. Catherall, A. Dorsival, V. Fedosseev, M. Georges, T. Giles, G. Le Scornet, U. Koester, J. Lettry, M. Lindroos, S. Marzari, S. Meunier, J. Parra-Lopez,

 

ISOLDE at URL: http://isolde.cern.ch

Consolidation projects at URL:

http://isolde-consolidation.web.cern.ch/isolde-consolidation/

PSMB meetings at URL:

http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html

 

1. HRS:

Tests have shown that the rise time of the 60kV power supply had risen from a normal 4ms to ~40ms. Further tests have revealed that this effect is due to the proton beam (3E+13 protons per pulse) hitting the thick converter. The secondary particles/neutrons produce a high ionization/current effect preventing the 60kV Astec power supply from reaching 60kV in the required time. Albeit reducing the HT to 52kV or reducing the proton beam intensity, this phenomenon is critical for the current experiment. Discussions with T. Fowler are for seen in attempt to try and improve this situation for future experiments. In the meantime, T. Giles suggests using the separator for Cs suppression instead of the converter.

The HRS logbook computer should be replaced as soon as possible.

 

2. GPS:

The target change was delayed by 24 hours due to a problem with the robot. The robot failed to place the new target on the target storage shelf before putting it on the front end. R. Catherall in collaboration with Staubli is investigating this problem.

J. Lettry has placed the previous target #233 behing the lead shielding doors.

     

3. LIS

V. Fedosseev has received the water analysis results that show that the water softening system does not work correctly. New resin and parts have been ordered for its repair.

 

4. Beam lines

After a recent intervention in the beam line power supply racks, M. Georges suggest that a synoptic for the identification of power supplies, interlocks and current use would be very useful. M. Lindroos reports that the LINUX synoptic being developed by the controls group could be used for this purpose.

S. Burger reports that the electronic card for the FNBS will be delivered in 2 to 3 weeks time.

 

 

5. Radiation Protection

The emergency handle of the HT access door has been forced in order to open the door. This is unacceptable especially since two type B (>2.5mSv/h) alarm levels were recorded this weekend. A TIS veto key, similar to the secondary zones, will be installed.

 

6. AOB

A simplified way of informing who is EIC and how to contact him would be most appreciated.

 

 

Absences

S. Marzari from

T. Giles from 13.09 to 07.10

T. Nilsson from 26.08 to 30.09 replaced by U. Koester

J. Schipper from 10.08 to 01.09

 

Next meeting:

Tuesday 3 September 2002 at 11.00 in the EP Conference Room A, building

304-1-001.